OpenStudy (anonymous):

Please help me understand finding the value of logarithms. a. log(2)16 b. log(5)25 c. log(4)64 d. log(5)625

OpenStudy (e.mccormick):

Are you looking for a way to aproximate it with a calculator or what?

OpenStudy (blurbendy):

These can be solved by simplifying things into exponents. i.e. log(2)16 = log(2)2^4 = 4

OpenStudy (anonymous):

To know how to solve it with and without the calculator.

OpenStudy (zzr0ck3r):

um this depends on the base of the log

OpenStudy (zzr0ck3r):

should that say\[\log_{2}(16)? \]

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

another way of saying that is 2^? = 16

OpenStudy (e.mccormick):

Because of what a log is: \(y=b^x\implies \log_b(y)=x\) Sometimes you can rearrange things to solve them.

OpenStudy (mertsj):

You need to know that the log is the exponent so when a problem asks you to find a log, it is asking you to find an exponent.

OpenStudy (zzr0ck3r):

5^? = 25

OpenStudy (anonymous):

Oh I get it now, thank-you

OpenStudy (zzr0ck3r):

4^? = 64

OpenStudy (mertsj):

Look at the first problem. It tells you that the base is 2 and the answer is 16. You are supposed to give the exponent that goes on the base 2 so that the answer will be 16.

OpenStudy (anonymous):

So 4, 5, 16, 125

OpenStudy (mertsj):

no.

OpenStudy (mertsj):

Look at the second problem. The base is 5. The answer is 25. What is the exponent that you would put on the base 5 to get an answer of 25?